~Include a normal curve/t distribution curve.

~Your final answer must be a concluding statement.


Suppose that the population standard deviation for a normally distributed standardized test of achievement is known to be 7.20. What would the standard error of the sample mean be if we were to draw a random sample of 16 test scores?


1
Expert's answer
2020-10-26T19:20:19-0400

standard error (SE) = sigma/"\sqrt{sample size}" = = 7.2/4 = 1.8